2x-1y+3z=5,3x+2y-1z=-2,1x+3y+4z=17
Answer by richwmiller(17219)

a 2-1+3=5,
b 3+2-1=-2,
c 1+3+4=17
a 2-1+3=5,
subtract a from c
a -1+4+1=12
b 3+2-1=-2,
c 1+3+4=17
add a and b
2+4+0=10
add a+c
0+7+5=29
2+4+0=10
etc
x = -1, y = 2, z = 3